# Information-Theoretic Thresholds for the Alignments of Partially Correlated Graphs

### v2

@@ 1 Introduction
The Erdős-Rényi random graph is the graph on $n$ vertices where each edge connects with probability $10<p<$ 1 independently. Let $\mathcal{G}(n,p)$ denote the distribution of Erdős-Rényi random graphs with $n$ vertices and edge connecting probability $p$.
@@
Given observations on $G_{1}$ and $G_{2}$ under the correlated Erdős-Rényi graphs model or correlated Gaussian Wigner model, the goal is to recover the latent vertex mapping $\pi$. Let $\mathrm{Supp}(\pi)$ denote the support set of $\pi$. To quantify the performance of an estimator $\hat{\pi}$, we consider the following two recovery criteria:
@@
    |$|\{v\in\mathrm{Supp}(\pi):\pi(v)=\hat{\pi}(v)\}|\geq\delta|\mathrm{Supp}(\pi)|.$ (1)
@@
    $\pi(v)=\hat{\pi}(v),\quad\forall~{}v\in\mathrm{Supp}(\pi).$ (2)
@@
  • Erdős-Rényi model, partial recovery. In the sparse regime where $np$ and $\rho$ are constant, partial recovery is impossible [GML21]. It is shown in [HM23] that $np(p\vee\rho)\gtrsim\log\left(1+\frac{\rho}{p}\right)\vee 1$ suffices for partial recovery, while $n\gtrsim d(p+\rho-p\rho\|p)\log n$ is necessary, where $d(p\|q)$ denotes the Kullback–Leibler (KL) divergence between Bernoulli distributions with mean $p$ and $q$, respectively. The recent work [WXY22] settled the sharp threshold for dense graphs with $\frac{p}{p\vee\rho}=n^{-o(1)}$ and the thresholds within a constant factor for sparse ones with $\frac{p}{p\vee\rho}=n^{-\Omega(1)}$. For the sparse case, a sharp threshold has been proven when $\frac{p}{p\vee\rho}=n^{-\alpha+o(1)}$ for $\alpha\in(0,1]$ in [DD23b].
@@
In this paper, we investigate the information-theoretic thresholds for recovering the set of correlated nodes $S^{*}$ and the mapping $\pi^{*}$. For notational simplicity, we also refer to the problem as recovering $\pi^{*}$ while keeping $S^{*}$ implicit as the domain of $\pi^{*}$. The success criteria are given by (1) and (2), where $\mathrm{Supp}(\pi)=S^{*}$. Owing to the potential inconsistency between the domain of $\pi^{*}$ and that of the estimator $\hat{\pi}:\hat{S}\mapsto V(G_{2})$, we define their overlap by
@@ 1.1 Main Results
In this subsection, we present the main results of the paper. We first introduce some notations for the presentation of the main theorems. Throughout the paper, we assume that $10<\rho<$ 1, $0<p\leq\frac{1}{2}$, and the cardinality $|S^{*}|=m$ is known. We denote the bivariate distribution of a pair of Bernoulli random variables with means $p_{1},p_{2}$, and correlation coefficient $\rho$ as $\mathrm{Bern}(p_{1},p_{2},\rho)$. Specifically, for $\mathrm{Bern}(p,p,\rho)$, we denote the following probability mass function:
@@
In the Erdős-Rényi model, we assume $p\geq\frac{1}{n}$, as partial recovery is otherwise impossible [WXY22]. Define
@@
Theorem 3 implies that in the Gaussian Wigner model, the thresholds for the number of correlated nodes $m$ are of the order $\frac{\log n}{\log\left(1/(1-\rho^{2})\right)}\vee 1$, beyond which exact recovery is possible and below which partial recovery is impossible. There is no gap between partial and exact recovery. Indeed, the estimation of $\pi^{*}$ can be divided into two steps: identifying the correlated nodes $(S^{*},T^{*})$ and recovering the mapping given $(S^{*},T^{*})$. Since the relative signal strength is characterized by the bounded correlation coefficient $\rho$, identifying the correlated nodes is always the major challenge, with no gap existing under the two recovery regimes.
@@ 1.2 Related Work
It is postulated in [HS17, Hop18, KWB19] that the framework of low-degree polynomial algorithms effectively demonstrates computation hardness of detecting and recovering latent structures, and it bears similarities to sum-of-square methods [HKP+17, Hop18]. Based on the conjecture on the hardness of low-degree polynomial algorithms, it is shown in [MWXY24] that there is no polynomial-time test or matching algorithm when the correlation coefficient satisfies $\rho^{2}\leq\frac{1}{\text{polylog}(n)}$. Furthermore, the recent work [DDL23] showed computation hardness for detection and exact recovery when $p=n^{-1+o(1)}$ and the correlation coefficient $\rho<\sqrt{\alpha}$, where $\alpha\approx 0.338$ is the Otter’s constant, suggesting that several polynomial algorithms may be essentially optimal.
The maximal overlap is a test statistic which aims to identify the mapping that maximizes the edge correlation between two graphs. It is known that finding the maximal overlap is an instance of quadratic assignment problem (QAP) [PRW94], which is NP-hard to solve or to approximate [MMS10]. There are many studies aiming to detect or recover latent structures based on the maximal overlap statistics [CK16, CK17, BCL+19, MX20, DMWX21, WXY22, WXY23, HM23]. Finally, we mention that the recent work [DDG24] approximated the maximal overlap within a constant factor in polynomial-time for sparse Erdős-Rényi graphs, and [DGH23] established a sharp transition on approximating problem on the performance of online algorithms for dense Erdős-Rényi graphs. In this paper, we do not address polynomial algorithms or the computational hardness on partial models and leave these topics for future work.
@@ 1.3 Notations
For any $n\in\mathbb{N}$, let $[n]\triangleq\{1,2,\cdots,n\}$. For any $a,b\in\mathbb{R}$, let $a\wedge b=\min\{a,b\}$ and $a\vee b=\max(a,b)$. We use standard asymptotic notation: for two positive sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}=O(b_{n})$ or $a_{n}\lesssim b_{n}$, if $a_{n}\leq Cb_{n}$ for some absolute constant $C$ and all $n$; $a_{n}=\Omega(b_{n})$ or $a_{n}\gtrsim b_{n}$, if $b_{n}=O(a_{n})$; $a_{n}=\Theta(b_{n})$ or $a_{n}\asymp b_{n}$, if $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$; $a_{n}=o(b_{n})$ or $b_{n}=\omega(a_{n})$, if $a_{n}/b_{n}\to 0$ as $n\to\infty$.
@@ 2 Correlated functional digraph
$V({\mathcal{H}}_{\pi}^{f})=V(G_{1}),\quad\beta_{e}(\mathcal{H}_{\pi}^{f})=\begin{cases}f\left(\beta_{e}(G_{1}),\beta_{\pi^{\mathsf{E}}(e)}(G_{2})\right),&\text{ if }e\in E(G_{1})\cap\binom{S}{2},\\ 0,&\text{ otherwise.}\end{cases}$
@@
It should be noted that both $\pi$ and $\pi^{*}$ are injective mappings under our model. After merging all pairs of nodes according to $\pi^{*}$, the degree of each vertex in the correlated functional digraph is at most two. Therefore, the connected components consist of paths and cycles, where a self-loop is understood as a cycle of length one. The connected components are illustrated in Figure 2. Let ${\mathcal{P}}$ and ${\mathcal{C}}$ denote the collections of subsets of ${\mathcal{E}}$ belonging to different connected paths and cycles, respectively. Note that the sets from ${\mathcal{P}}$ and ${\mathcal{C}}$ are disjoint. Consequently,
@@
For simplicity, we write $\kappa_{\ell}^{\mathsf{C}}(t)=\kappa_{\ell}^{{\mathsf{C}},f}(t)$ and $\kappa_{\ell}^{\mathsf{P}}(t)=\kappa_{\ell}^{{\mathsf{P}},f}(t)$ when the function $f$ is specified. Lower-order cumulants can be calculated directly. See more details in Appendix B.1. However, it is crucial to establish upper bounds for higher-order cumulants in terms of the lower-order ones. To this end, we introduce the following lemma.
@@
The proof of Lemma 1 is deferred to Appendix B.1. The special case where both $\pi$ and $\pi^{*}$ are bijective has been studied in [WXY22, DD23b, HM23], where the correlation relationships can be characterized by the permutation $(\pi^{*})^{-1}\circ\pi$. In this case, the connected components of the functional digraph of permutations are all cycles. In contrast, in our setting, the domains and ranges of $\pi$ and $\pi^{*}$ may differ, requiring us to deal with the intricate correlations among edges involving both cycles and paths, as addressed by Lemma 1.
@@
$f(x,y)=-\frac{\rho^{2}}{2}\left(x^{2}+y^{2}\right)+\rho xy.$ (9)
@@ 3 Recovery by maximizing total similarity score
$\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})>\max_{\pi:d(\pi,\pi^{*})\geq d_{0}}\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})=\max_{k\geq d_{0}}\max_{\pi:d(\pi,\pi^{*})=k}\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})$
with high probability, where the thresholds $d_{0}=1$ and $(1-\delta)m$ correspond to exact and partial recoveries, respectively. In the following, we outline a general recipe to derive an upper bound for the failure event $\{\max_{\pi\in\mathcal{T}_{k}}\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})\geq\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})\}$ for a fixed $k$, where $\mathcal{T}_{k}\subseteq\mathcal{S}_{n,m}$ denotes the set of injections $\pi$ such that $d(\pi,\pi^{*})=k$. The error probabilities in the main theorems are then obtained by summing over the corresponding range of $k$.
For any $\pi\in\mathcal{T}_{k}$, by definition, there exists a set of correctly matched vertices of cardinality $m-k$—corresponding to the self-loops in the correlated functional digraph of $\pi$ over the vertices—denoted by $F_{\pi}\triangleq\{v\in S^{*}\cap S:\pi^{*}(v)=\pi(v)\}$ with $|F_{\pi}|=m-k$. The induced subgraphs of ${\mathcal{H}}_{\pi}^{f}$ and ${\mathcal{H}}_{\pi^{*}}^{f}$ over $F_{\pi}$ are identical. Consequently,
@@
For a fixed $\pi\in\mathcal{T}_{k}$, the random variable $\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])$ represents the total weight across $N_{k}\triangleq\binom{m}{2}-\binom{m-k}{2}=mk(1-\frac{k+1}{2m})$ pairs of vertices. Additionally, $F_{\pi}$ is a subset of $S^{*}$ of cardinality $m-k$, and the total number of possible configurations for $F_{\pi}$ is at most $\binom{m}{m-k}=\binom{m}{k}$. Therefore,
@@
It remains to upper bound the total number of $\pi$ in $\mathcal{T}_{k}$. To do so, we first choose $m-k$ elements from the domain of $\pi^{*}$ and map them to the same value as $\pi^{*}$. Then, the remaining domain and range, each of size $k$, are matched arbitrarily. This yields:

### v3

@@ 1 Introduction
The Erdős-Rényi random graph is the graph on $n$ vertices where each edge connects with probability $0<p<1$ independently. Let $\mathcal{G}(n,p)$ denote the distribution of Erdős-Rényi random graphs with $n$ vertices and edge connecting probability $p$.
@@
Given observations on $G_{1}$ and $G_{2}$ under the correlated Erdős-Rényi graphs model or correlated Gaussian Wigner model, the goal is to recover the latent vertex mapping $\pi$. To quantify the performance of an estimator $\hat{\pi}$, we consider the following two recovery criteria:
@@
    $|\{v\in{V(G_{1})}:\pi(v)=\hat{\pi}(v)\}|\geq\delta|{V(G_{1})}|.$ (1)
@@
    $\pi(v)=\hat{\pi}(v),\quad\forall~v\in{V(G_{1})}.$ (2)
@@
  • Erdős-Rényi model, partial recovery. In the sparse regime where $np$ and $\rho$ are constant, partial recovery is impossible when $n(p^{2}+\rho p(1-p))\leq 1$ [GML21, WXY22]. It is shown in [HM23] that $np(p\vee\rho)\gtrsim\log\left(1+\frac{\rho}{p}\right)\vee 1$ suffices for partial recovery, while $n\gtrsim d_{\mathrm{KL}}(p+\rho-p\rho\|p)\log n$ is necessary, where $d_{\mathrm{KL}}(p\|q)$ denotes the Kullback–Leibler (KL) divergence between Bernoulli distributions with mean $p$ and $q$, respectively. The recent work [WXY22] settled the sharp threshold for dense graphs with $\frac{p}{p\vee\rho}=n^{-o(1)}$ and the thresholds within a constant factor for sparse ones with $\frac{p}{p\vee\rho}=n^{-\Omega(1)}$. For the sparse case, a sharp threshold has been proven when $\frac{p}{p\vee\rho}=n^{-\alpha+o(1)}$ for $\alpha\in(0,1]$ in [DD23b].
@@
In this paper, we investigate the information-theoretic thresholds for recovering the set of correlated nodes $S^{*}$ and the mapping $\pi^{*}$. For notational simplicity, we also refer to the problem as recovering $\pi^{*}$ while keeping $S^{*}$ implicit as the domain of $\pi^{*}$. The success criteria in the fully correlated graph models are given by (1) and (2). In the partially correlated graph models, owing to the potential inconsistency between the domain of $\pi^{*}$ and that of the estimator $\hat{\pi}:\hat{S}\mapsto V(G_{2})$, we define their overlap by
@@ 1.1 Main Results
In this subsection, we present the main results of the paper. We first introduce some notations for the presentation of the main theorems. Throughout the paper, we assume that $0<\rho<1$, $0<p\leq\frac{1}{2}$, and the cardinality $|S^{*}|=m$ is known. We denote the bivariate distribution of a pair of Bernoulli random variables with means $p_{1},p_{2}$, and correlation coefficient $\rho$ as $\mathrm{Bern}(p_{1},p_{2},\rho)$. Specifically, for $\mathrm{Bern}(p,p,\rho)$, we denote the following probability mass function:
@@
In the Erdős-Rényi model, we assume $p\geq\frac{1}{n}$, as partial recovery is otherwise impossible [GML21, WXY22]. Define
@@
Theorem 3 implies that in the Gaussian Wigner model, the thresholds for the number of correlated nodes $m$ are of the order $\frac{\log n}{\log\left(1/(1-\rho^{2})\right)}\vee 1$, beyond which exact recovery is possible and below which partial recovery is impossible. There is no gap between partial and exact recovery.
@@
1.2 Interpretation of Recovery Thresholds
Estimating $\pi^{*}$ under the partially correlated graph model conceptually consists of two subproblems: 1) recovery of the support sets $S^{*}$ and $T^{*}$; and 2) recovery of the matching between the two sets. The existing literature mostly focuses on the latter, yielding necessary conditions under Erdős-Rényi and Gaussian Wigner models, respectively. In our results, new thresholds $\frac{\log n}{p^{2}\phi(\gamma)}$ and $\frac{\log n}{\log(1/(1-\rho^{2}))}$ reflect the complexity of the support recovery problem. We provide more discussions on the support recovery problem in Appendix A.
Those new thresholds have a natural information-theoretical interpretation. The entropy of $\pi^{*}$ (or the support sets $S^{*}$ and $T^{*}$) is on the order of $m\log n$. On the other hand, the mutual information between $\pi^{*}$ and the observed pair of graphs $(G_{1},G_{2})$ can be upper bounded as follows:
  • Erdős-Rényi model: $I(\pi^{*};G_{1},G_{2})\leq 25\binom{m}{2}p^{2}\phi(\gamma)$,
  • Gaussian Wigner model: $I(\pi^{*};G_{1},G_{2})\leq\frac{1}{2}\binom{m}{2}\log(\frac{1}{1-\rho^{2}})$,
where $p^{2}\phi(\gamma)$ and $\log(\frac{1}{1-\rho^{2}})$ arise from the KL divergence between a pair of correlated edges and independent edges (see Lemma 5). Combining those yields the necessary conditions on the exact recovery by
$I(\pi^{*};G_{1},G_{2})\gtrsim m\log n.$
For partial recovery, similar arguments can be applied to the metric entropy of $\pi^{*}$ using standard volume argument. Then the necessary conditions on recovery thresholds follow from Fano’s method.
The key quantities $p^{2}\phi(\gamma)$ and $\log(\frac{1}{1-\rho^{2}})$ also have an intuitive large deviation interpretation from the upper bounds. In our estimator, an edge pair with $e_{1}\in E(G_{1})$ and $e_{2}\in E(G_{2})$ contributes $f(\beta_{e_{1}}(G_{1}),\beta_{e_{2}}(G_{2}))$ to the total similarity score, where $f$ is a prescribed similarity function. An incorrect matching involves two independent edges, while a correct matching consists of correlated edges with a higher expected score. For the Erdős-Rényi model, we use $f(x,y)=xy$ and the rate function analysis of Bernoulli distributions yields $d_{\mathrm{KL}}(p_{11}\|p^{2})\geq p^{2}\phi(\gamma)$ [^1] This is also known as Bennett inequality.; for the Gaussian Wigner model, the quantity $\log(\frac{1}{1-\rho^{2}})$ can be obtained by combining the rate function analyses of Gaussian distributions with $f(x,y)=xy$ and $f(x,y)=-\frac{1}{2}(x-y)^{2}$. However, classical large deviation theory is not directly applicable due to the correlations among the incorrectly matched edges. We address this issue by carefully analyzing the correlation structures in Section 2.
To illustrate the recovery thresholds, consider $p=n^{-a_{1}}$ and $\rho=n^{-a_{2}}$ for constants $a_{1},a_{2}\in(0,1)$ in the Erdős-Rényi model. The exact recovery and partial recovery thresholds are given by
Exact Recovery: $\begin{cases}n^{a_{1}+a_{2}}\log n,&a_{1}>a_{2}\\ n^{a_{1}+a_{2}}\log n,&a_{1}=a_{2}\\ n^{2a_{2}}\log n,&a_{1}<a_{2}\end{cases},\quad\text{Partial Recovery: }\begin{cases}n^{a_{1}+a_{2}},&a_{1}>a_{2}\\ n^{a_{1}+a_{2}}\log n,&a_{1}=a_{2}\\ n^{2a_{2}}\log n,&a_{1}<a_{2}\end{cases}.$
There is a transition in the recovery thresholds from $n^{a_{1}+a_{2}}\log n$ to $n^{2a_{2}}\log n$ as the relative signal strength $\gamma=\frac{\rho(1-p)}{p}$ changes from the weak signal regime to the strong signal regime. Let $m=n^{a_{3}}$. The critical threshold of the exponent $a_{3}$ as a function of parameters $a_{1}$ and $a_{2}$ is illustrated in the contour plot in Figure 1. In this phase diagram, we focus on the polynomial dependence on $n$ and ignore logarithmic factors. The red line with $m=n$ was established in prior work [WXY22]. Our contribution is to determine the critical recovery thresholds on $m$ in the green region for all configurations of $p$ and $\rho$.
[…]
Figure 1: Phase diagram for recovery thresholds with $p=n^{-a_{1}}$, $\rho=n^{-a_{2}}$, and $m=n^{a_{3}}$.
@@ 1.3 Related Work
Subgraphs isomorphism problem.
In our partially correlated Erdős-Rényi model, there are two latent subgraphs of order $m$ with correlated edges. Specifically, when $\rho=1$, the two subgraphs are isomorphic. Intuitively, a necessary condition for a successful alignment is that $m$ exceeds the order the largest induced isomorphic subgraph between the independent parts. It is recently shown in [CD23, SWZ25] that the size of maximal isomorphic subgraphs between two independent Erdős-Rényi graphs $\mathcal{G}(n,p)$ is $\Theta(\log n)$ when $p$ is a constant. In this regime, our threshold $m\geq C\log n$ for some constant $C$ aligns with such results. When $p=p_{n}\to 0$, the order of the largest isomorphic subgraphs between two independent Erdős-Rényi graphs remains open [SWZ25].
@@
It is postulated in [HS17, Hop18, KWB19] that the framework of low-degree polynomial algorithms effectively demonstrates computation hardness of detecting and recovering latent structures, and it bears similarities to sum-of-square methods [HKP+17, Hop18]. Based on the conjecture on the hardness of low-degree polynomial algorithms, it is shown in [MWXY24] that there is no polynomial-time test or matching algorithm when the correlation coefficient satisfies $\rho^{2}\leq\frac{1}{\text{polylog}(n)}$. Furthermore, the recent work [DDL23] showed computation hardness for detection and exact recovery when $p=n^{-1+o(1)}$ and the correlation coefficient $\rho<\sqrt{\alpha}$, where $\alpha\approx 0.338$ is the Otter’s constant, suggesting that several polynomial algorithms may be essentially optimal. In this paper, we do not address polynomial algorithms or the computational hardness on partial models and leave these topics for future work.
@@ 1.4 Notations
For any $n\in\mathbb{N}$, let $n]\triangleq\{1,2,\cdots,n\}$. For any $a,b\in\mathbb{R}$, let $a\wedge b=\min\{a,b\}$ and $a\vee b=\max(a,b)$. We use standard asymptotic notation: for two positive sequences $\{a_{n}\}$ and $\{b_{n}\}$, we write $a_{n}=O(b_{n})$ or $a_{n}\lesssim b_{n}$, if $a_{n}\leq Cb_{n}$ for some absolute constant $C$ and all $n$; $a_{n}=\Omega(b_{n})$ or $a_{n}\gtrsim b_{n}$, if $b_{n}=O(a_{n})$; $a_{n}=\Theta(b_{n})$ or $a_{n}\asymp b_{n}$, if $a_{n}=O(b_{n})$ and $a_{n}=\Omega(b_{n})$; $a_{n}=o(b_{n})$ or $b_{n}=\omega(a_{n})$, if $a_{n}/b_{n}\to 0$ as $n\to\infty$.
@@ 2 Correlated functional digraph
[…]
@@
It should be noted that both $\pi$ and $\pi^{*}$ are injective mappings under our model. After merging all pairs of nodes according to $\pi^{*}$, the degree of each vertex in the correlated functional digraph is at most two. Therefore, the connected components consist of paths and cycles, where a self-loop is understood as a cycle of length one. The connected components are illustrated in Figure 3. Following prior works on graph matching such as [RS21] and [YC23], we lift the vertex mapping $\pi$ to an edge mapping $\pi^{{\mathsf{E}}}$. When the support sets are known, the lifted permutations can be decomposed into cycles as independent components. This viewpoint is consistent with the standard cycle-path decomposition that has been fruitfully used in graph alignment analysis. For instance, [CKMP20] explicitly decomposed the lifted matching into cycles and paths in their partial-recovery analysis. [GRS22] employed a lifted-object perspective that accommodates path and cycle components when analyzing correlated stochastic block models. [OGE16] studied de-anonymization under community structure, where analogous decompositions are handled in a more general fashion. Our contribution is to tailor this decomposition to the partially correlated setting with unknown support, and to exploit the resulting factorization in our threshold analysis.
Let $\mathcal{P}$} and $\mathcal{C}$} denote the collections of subsets of $\mathcal{E}$} belonging to different connected paths and cycles, respectively. Note that the sets from $\mathcal{P}$} and $\mathcal{C}$} are disjoint. Consequently,
@@
For simplicity, we write $\kappa_{\ell}^{\mathsf{C}}(t)=\kappa_{\ell}^{{\mathsf{C}},f}(t)$ and $\kappa_{\ell}^{\mathsf{P}}(t)=\kappa_{\ell}^{{\mathsf{P}},f}(t)$ when the function $f$ is specified. Lower-order cumulants can be calculated directly. See more details in Appendix C.1. However, it is crucial to establish upper bounds for higher-order cumulants in terms of the lower-order ones. To this end, we introduce the following lemma.
@@
The proof of Lemma 1 is deferred to Appendix C.1. The special case where both $\pi$ and $\pi^{*}$ are bijective has been studied in [WXY22, DD23b, HM23], where the correlation relationships can be characterized by the permutation $(\pi^{*})^{-1}\circ\pi$. In this case, the connected components of the functional digraph of permutations are all cycles. In contrast, in our setting, the domains and ranges of $\pi$ and $\pi^{*}$ may differ, requiring us to deal with the intricate correlations among edges involving both cycles and paths, as addressed by Lemma 1.
Remark 2.
For the Erdős-Rényi model, when $f(x,y)=xy$, the estimator (6) is equivalent to computing the maximal edge overlap over all possible injective mappings. The recent work [DDG24] approximated the maximal edge overlaps within a constant factor in polynomial-time for sparse Erdős-Rényi graphs, and [DGH23] established a sharp transition on approximating problem on the performance of online algorithms for dense Erdős-Rényi graphs. Assume $p=n^{-a_{1}}$, $\rho=n^{-a_{2}}$, and $m=n^{a_{3}}$. It is shown in [DDG24] that the maximal edge overlap between two independent Erdős-Rényi graphs $\mathcal{G}(m,p)$ is $\frac{m}{2\alpha-1}$ when $p=m^{-\alpha}$ for $\alpha\in(1/2,1)$. In contrast, in the partially correlated model, the maximal edge overlap is approximately $\binom{m}{2}p_{11}\asymp n^{2a_{3}-a_{1}-a_{1}\wedge a_{2}}.$ Under the condition for successful recovery, we have $a_{3}\geq a_{2}+a_{1}\vee a_{2}\geq a_{1}+a_{1}\wedge a_{2}$, and it follows that $\binom{m}{2}p_{11}\gtrsim n^{a_{3}}\asymp\frac{m}{2\alpha-1}.$ Therefore, the maximal edge overlap in the partially correlated model exceeds that in the independent case.
@@
$f_{\mathrm{MLE}}(x,y)=-\frac{\rho}{2}\left(x^{2}+y^{2}\right)+xy.$ (9)
Specifically,
  • when $1-\rho=\Omega(1)$, $f_{\mathrm{MLE}}(x,y)\asymp xy$;
  • when $1-\rho=o(1)$, $f_{\mathrm{MLE}}(x,y)\asymp-\frac{1}{2}(x-y)^{2}$.
@@ 3 Recovery by maximizing total similarity score
$\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})>\max_{\pi:d(\pi,\pi^{*})\geq d_{0}}\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})=\max_{k\geq d_{0}}\max_{\pi:d(\pi,\pi^{*})=k}\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})$ (10)
with high probability, where the thresholds $d_{0}=1$ and $(1-\delta)m$ correspond to exact and partial recoveries, respectively. Recall that the weight in $\mathcal{H}_{\pi}^{f}$ represents the similarity score under injection $\pi$, and our estimator $\hat{\pi}$ maximizes this similarity score. To achieve partial or exact recovery, it suffices to show that the total similarity score under the true mapping $\pi^{*}$ is larger as in (10).
In the following, we outline a general recipe to derive an upper bound for the failure event $\{\max_{\pi\in\mathcal{T}_{k}}\mathsf{{e}}({\mathcal{H}}_{\pi}^{f})\geq\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})\}$ for a fixed $k$, where $\mathcal{T}_{k}\subseteq\mathcal{S}_{n,m}$ denotes the set of injections $\pi$ such that $d(\pi,\pi^{*})=k$. The error probabilities in the main theorems are then obtained by summing over the corresponding range of $k$.
For any $\pi\in\mathcal{T}_{k}$, by definition, there exists a set of correctly matched vertices of cardinality mkm-k—corresponding to the self-loops in the correlated functional digraph of $\pi$ over the vertices—denoted by $F_{\pi}\triangleq\{v\in S^{*}\cap S:\pi^{*}(v)=\pi(v)\}$ with $|F_{\pi}|=m-k$. By the definition of $F_{\pi}$, the two induced subgraphs ${\mathcal{H}}_{\pi}^{f}[F_{\pi}]$ and ${\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}]$ are identical, and so are the corresponding edge summations $\mathsf{{e}}(\mathcal{H}_{\pi}^{f}[F_{\pi}])=\mathsf{{e}}(\mathcal{H}_{\pi^{*}}^{f}[F_{\pi}])$. Consequently,
@@
For a fixed $\pi\in\mathcal{T}_{k}$, the random variable $\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])$ represents the total weight across $N_{k}\triangleq\binom{m}{2}-\binom{m-k}{2}=mk(1-\frac{k+1}{2m})$ pairs of vertices. Additionally, $F_{\pi}$ is a subset of $S^{*}$ of cardinality mkm-k, and the total number of possible configurations for $F_{\pi}$ is at most $\binom{m}{m-k}=\binom{m}{k}$. Therefore,
@@
It remains to upper bound the total number of $\pi$ in $\mathcal{T}_{k}$. To do so, we first choose mkm-k elements from the domain of $\pi^{*}$ and map them to the same value as $\pi^{*}$. Then, the remaining domain and range, each of size $k$, are matched arbitrarily. This yields: